# TimeSpot: Benchmarking Geo-Temporal Understanding in Vision–Language Models in Real-World Settings

### 04 Mar 2026 v1

@@ 2 Preliminaries and Related Work
| Countries | 82 unique; top: USA (196), Russia (97), Japan (67), Italy (65), China (58) | |
@@
Practical Significance of Temporal Reasoning. Accurate temporal inference from visual input is essential for real-world systems in which decisions depend not only on location, but on when an observation occurred relative to environmental, social, and physical dynamics. Accurate temporal understanding underpins a wide range of applications, from surveillance and visual event analysis to embodied agents, traffic coordination, and disaster management in real-world settings. In applications such as disaster management, traffic management, and environmental monitoring, the same location can imply fundamentally different risks, system states, or actions depending on season, daylight phase, or time of day (tehrani2026disasterinsightmultimodalbenchmarkfunctionaware; liao2025workzoneschallengevlm). From a world-modeling perspective, temporal reasoning underpins coherent internal representations that support prediction, planning, counterfactual reasoning, and physical plausibility, rather than static scene recognition (wu2025chronoedittemporalreasoningimage). Prior work shows that models incorporating explicit temporal reasoning outperform spatial-only systems in time-sensitive tasks such as traffic forecasting and mobility analysis, where errors of hours can translate into operational failure (Zhou2025awfa). Moreover, temporal misalignment has been identified as a key source of brittleness in deployed vision and multimodal systems, leading to physically implausible behavior and unsafe decisions in embodied and autonomous settings despite high spatial accuracy (matta2025waydoestimeflow; li2026finegrained). Our findings demonstrate that current VLMs frequently fail at this temporal grounding even when spatial predictions appear correct, highlighting geo-temporal reasoning as a prerequisite for reliable real-world deployment.
@@ 3.1 Benchmark Construction
All samples undergo two-stage verification. Annotators first collect auxiliary geographic context and cross-check programmatic labels against visual cues such as shadows, illumination, vegetation, and weather. Senior annotators then adjudicate flagged cases, including twilight ambiguity and artificial lighting, by validating metadata against ephemeris calculations and scene consistency.
@@ 4 Experiments and Evaluation
Evaluation. Evaluation is interpretable and verifiable by design. We report categorical accuracy for continent, country, climate zone, and environment type; local-time accuracy within 1 hour and mean absolute error (MAE, minutes); coordinate MAE (degrees) and mean geodesic distance (MD, km) based on great-circle distance (tian2017cross; vo2016localizing; arandjelovic2016netvlad). To assess trustworthiness, we include explicit cross-field geo-temporal consistency diagnostics, calibration metrics such as expected calibration error and risk–coverage curves, and robustness analyses via stratification across continents, climate zones, and environment types, as well as hemisphere-flip tests and hard out-of-distribution splits (zhu2021vigor; astruc2024openstreetview; hou2024global; huang2024cv). This evaluation isolates joint geo-temporal reasoning capabilities that generic VLM benchmarks (radford2021learning; kim2021vilt; li2022blip) and remote-sensing-focused suites such as GEOBench-VLM (danish2024geobench) do not explicitly test, namely calibrated and physically consistent inference of both time and place from subtle ground-level visual evidence. Formal metric definitions and scoring thresholds are provided in D.2 and prompts in E.
Improving Visual Reasoning via Supervised Fine-Tuning (SFT). As a diagnostic intervention, we study whether supervised fine-tuning can improve geo-temporal reasoning on TimeSpot. We fine-tune Qwen-VL2.5-3B-Instruct on 40% of the dataset using stratified sampling and evaluate on the remaining 60% (H.1). SFT yields consistent reductions in spatial and temporal errors, indicating improved grounding in relational and geometric cues, but does not eliminate core failure modes. These results suggest that while explicit supervision can improve generalization, but it is insufficient on its own to induce robust geo-temporal reasoning.
@@ 5.4 Performance Improvement with SFT
Figure 4 reports supervised fine-tuning results on TimeSpot (details in H.1). Fine-tuning Qwen2.5-VL-3B-Instruct on a stratified 40% subset yields consistent improvements in country accuracy, rising from 14.20% to 19.24% by epoch 4, demonstrating that SFT effectively strengthens categorical geo-semantic prediction. In contrast, time accuracy improves early (20.27% to 23.37%) but exhibits noticeable fluctuations across later epochs, indicating less stable learning of fine-grained temporal cues. This asymmetry suggests that SFT more reliably reinforces discrete place identity than continuous temporal inference from illumination, shadow geometry, and solar dynamics. Overall, SFT provides a meaningful and reliable boost to geo-temporal performance on TimeSpot, while highlighting the need for complementary approaches that explicitly target temporal grounding and physical reasoning beyond fine-tuning alone.
@@ 6 Concluding Remarks
Geo-temporal reasoning remains a major challenge for vision–language models in unconstrained, real-world images. We show that joint geo-temporal reasoning remains a core unresolved challenge for vision–language models operating on unconstrained real-world images. We introduce TimeSpot, a benchmark of 1,455 ground-level images from 80 countries that requires explicit prediction of when and where a scene occurs through structured temporal and geographic attributes. Our analysis exposes failure modes that are invisible to spatial-only or retrieval-based benchmarks. Empirically, even the strongest models exhibit large geodesic errors and low time-of-day accuracy, with frequent violations of solar, hemispheric, and daylight consistency. These errors reveal over-reliance on coarse spatial priors and weak grounding in physical cues such as illumination, shadows, and seasonal dynamics. Importantly, temporal reasoning is not ancillary but essential for coherent world modeling, physically plausible inference, and downstream decision-making in real-world systems. Future works should prioritize explicit geo-temporal inductive biases, constraint-aware reasoning, and supervision that couples spatial context with temporal physics and dynamics.
@@ H.1.2 Supervised Fine-Tuning (SFT) Performance
Figure Figure 4 summarize the effect of supervised fine-tuning on TimeSpot across training epochs. Starting from the pretrained baseline (epoch 0), SFT yields consistent improvements in country accuracy, rising from 14.20% to 19.24% by epoch 4, with performance remaining stable thereafter. This trend indicates that supervised adaptation effectively strengthens categorical geo-semantic discrimination, enabling the model to better associate visual cues with country-level identity.
@@ I.1 Limitations of Current Vision–Language Models
Limits of supervised fine-tuning. Supervised fine-tuning provides consistent gains for categorical geo-semantic prediction but yields smaller and less stable improvements for continuous temporal inference. SFT tends to reinforce frequent patterns aligned with structured labels, while struggling to induce robust reasoning over solar geometry and illumination physics. This suggests that standard fine-tuning objectives are not well matched to the demands of physically grounded temporal reasoning.
@@ I.2 Future Research Directions
Beyond fine-tuning: architectural and training innovations. While SFT improves baseline performance, our results indicate that deeper progress will require architectural changes and training paradigms that explicitly model time as part of world representation. Promising directions include hybrid perception–physics models, modular temporal reasoning components, and multi-task training that couples spatial perception with temporal prediction.
@@
Appendix J Examples from TimeSpot
@@
Appendix K Examples of JSON record with Response Analysis

### 25 May 2026 v2

@@ 2 Preliminaries and Related Work
| Countries | 80 unique; top: USA (196), Russia (97), Japan (67), Italy (65), China (58) | |
@@
Practical Significance of Temporal Reasoning. Beyond cognitive interest, temporal inference directly affects systems where the same location implies different risks, operational states, or decisions depending on season, daylight phase, or time of day (tehrani2026disasterinsightmultimodalbenchmarkfunctionaware; liao2025workzoneschallengevlm). Temporal misalignment has been identified as a source of physically implausible behavior in deployed vision systems even when spatial predictions are high (matta2025waydoestimeflow). Our findings confirm this: VLMs frequently fail on temporal grounding even when spatial predictions are correct, reinforcing that geo-temporal reasoning is a prerequisite for trustworthy real-world deployment.
@@ 3.1 Benchmark Construction
All samples undergo two-stage verification. Annotators first collect auxiliary geographic context and cross-check programmatic labels against visual cues such as shadows, illumination, vegetation, and weather. Senior annotators then adjudicate flagged cases, including twilight ambiguity and artificial lighting, by validating metadata against ephemeris calculations and scene consistency. The annotation team comprised 3 primary annotators and 2 senior annotators, all engineering graduates with geospatial domain experience; no crowdsourcing platforms were used. Primary annotation required approximately 576 hours across 6 weeks; senior adjudication of edge cases required an additional 30–40 hours.
@@
| Human Baselines | ||||||||||||
| Average (Undergrad) | 80.89 | 45.98 | 67.96 | 75.71 | 12.89 | 33.06 | 2800.49 | 68.89 | 28.39 | 41.92 | 2:41 | 77.89 |
| Average (Expert) | 94.06 | 67.89 | 86.39 | 87.39 | 5.16 | 12.22 | 1040.42 | 86.56 | 46.06 | 57.89 | 1:36 | 92.22 |
@@ 4 Experiments and Evaluation
Evaluation. All proprietary model evaluations used the OpenRouter API; total inference cost was approximately 1,450 USD, with an estimated 400–500 million tokens across all runs (model pricing: 2–15 USD per million tokens depending on capability tier). We report categorical accuracy for continent, country, climate zone, and environment type; local-time accuracy within 1 hour and mean absolute error (MAE, minutes); coordinate MAE (degrees) and mean geodesic distance (MD, km) based on great-circle distance (tian2017cross; vo2016localizing; arandjelovic2016netvlad). To assess trustworthiness, we include explicit cross-field geo-temporal consistency diagnostics, calibration metrics such as expected calibration error and risk–coverage curves, and robustness analyses via stratification across continents, climate zones, and environment types, as well as hemisphere-flip tests and hard out-of-distribution splits (zhu2021vigor; astruc2024openstreetview; hou2024global; huang2024cv). This evaluation isolates joint geo-temporal reasoning capabilities that generic VLM benchmarks (radford2021learning; kim2021vilt; li2022blip) and remote-sensing-focused suites such as GEOBench-VLM (danish2024geobench) do not explicitly test, namely calibrated and physically consistent inference of both time and place from subtle ground-level visual evidence. Formal metric definitions and scoring thresholds are provided in D.2 and prompts in E.
Improving Visual Reasoning via Supervised Fine-Tuning (SFT). As a diagnostic intervention, we investigate whether supervised fine-tuning improves geo-temporal reasoning on TimeSpot. We fine-tune Qwen-VL2.5-3B-Instruct on country and time prediction tasks using a stratified 40% training split and evaluate on the remaining 60% (H.1). The reduced training split is designed to test generalization rather than memorization of geographic distributions. While SFT improves both country and time prediction over the zero-shot baseline, cross-task evaluation reveals interference between spatial and temporal objectives under single-task LoRA adaptation. Joint SFT (H.2) partially mitigates this trade-off but remains limited by competing gradient signals from illumination-invariant and illumination-sensitive cues. These findings motivate future RL-based or constraint-aware training strategies better aligned with physically grounded geo-temporal reasoning (chen2025sftrl).
@@ 5.1 Performance Analysis
Human performance baseline. To establish empirical difficulty bounds for the benchmark tasks, we conducted a controlled human evaluation involving two groups given only the image (no metadata): average participants (undergraduates) and domain experts (geology and geography graduates). Expert participants achieved a Time MAE of 1:36 and absolute Time Accuracy (within 1 hour) of 57.89%, Season accuracy of 86.56%, Country accuracy of 67.89%, Climate accuracy of 86.39%, and Daylight Phase accuracy of 92.22%. Average participants scored lower across all fields (Time MAE: 2:41, Season: 68.89%, Country: 45.98%). These results confirm that (i) the 1-hour time accuracy threshold is empirically achievable by domain experts reasoning from physical cues alone; (ii) VLMs substantially lag human experts on all temporal fields (best VLM Time MAE: 3:36 vs. expert: 1:36); and (iii) the best VLMs match or exceed both human groups on coarse spatial memorization (e.g., Gemini-2.5-Flash country accuracy: 77.25% vs. expert: 67.89%), while lacking the continuous 4D physical-world understanding that underlies expert temporal inference. Full human baseline results are reported in Table 25.
@@ 5.4 Performance Improvement with SFT
Figure 4 reports supervised fine-tuning results on TimeSpot (details in H.1). Fine-tuning Qwen2.5-VL-3B-Instruct on a stratified 40% subset improves both country accuracy (14.20% 19.24% by epoch 4) and time accuracy (20.27% 24.79% by epoch 5), showing that SFT strengthens geo-semantic prediction and temporal grounding. However, country accuracy saturates early, while time accuracy fluctuates across epochs, indicating less stable learning of fine-grained temporal cues. Cross-task evaluation reveals feature interference: country-tuned models reduce time accuracy (22.06% 21.78%), while time-tuned models reduce country accuracy (13.47% 12.98%). This reflects competing visual demands: country prediction relies on illumination-invariant structural cues, whereas time prediction depends on illumination-sensitive features such as shadows and sky luminance. Joint SFT partially mitigates this trade-off (country: 14.23% 15.72%; time: 20.27% 22.36%), but remains below single-task peaks, suggesting gradient conflict under shared LoRA parameters. These findings motivate training methods that better separate or balance spatial and temporal gradient signals (schulzebuschoff2025finetuning; binz2025foundation; chen2025sftrl).
@@ 6 Concluding Remarks
Geo-temporal reasoning remains a major challenge for vision–language models in unconstrained, real-world images. In this work, we show that joint geo-temporal reasoning remains a core unresolved challenge for vision–language models operating on unconstrained real-world images. We introduce TimeSpot, a benchmark of 1,455 ground-level images from 80 countries that requires explicit prediction of when and where a scene occurs through structured temporal and geographic attributes. Our analysis exposes failure modes that are invisible to spatial-only or retrieval-based benchmarks. Empirically, even the strongest models exhibit large geodesic errors and low time-of-day accuracy, with frequent violations of solar, hemispheric, and daylight consistency. These errors reveal over-reliance on coarse spatial priors and weak grounding in physical cues such as illumination, shadows, and seasonal dynamics. Importantly, temporal reasoning is not ancillary but essential for coherent world modeling, physically plausible inference, and downstream decision-making in real-world systems. Future works should prioritize explicit geo-temporal inductive biases, constraint-aware reasoning, and supervision that couples spatial context with temporal physics and dynamics.
@@
Conflict of Interest Disclosure
The authors declare that they have no financial or other substantive conflicts of interest that could reasonably be perceived to influence this work.
@@ H.1.2 Supervised Fine-Tuning (SFT) Performance
Figure 4 summarizes the effect of supervised fine-tuning on TimeSpot across training epochs. Starting from the pretrained baseline (epoch 0), SFT yields consistent improvements in country accuracy, rising from 14.20% to 19.24% by epoch 4, with performance remaining stable thereafter. This trend indicates that supervised adaptation effectively strengthens categorical geo-semantic discrimination, enabling the model to better associate visual cues with country-level identity.
@@
H.1.3 Cross-task Evaluation
To assess whether single-task SFT generalizes across domains, we evaluate each fine-tuned checkpoint on the complementary task. As shown in Table 24, fine-tuning on country prediction reduces time accuracy from 22.06% to 21.78%, while fine-tuning on time prediction reduces country accuracy from 13.47% to 12.98%. This interference reflects that country prediction requires invariance to illumination (for lighting-robust place identity), while time prediction requires sensitivity to illumination (for shadow and sky state inference). Under shared LoRA parameters, these competing gradient directions reduce task-specific specialization.
| Fine-tuned on | Country Ac. (%) | Time Ac. (%) |
| Baseline (no SFT) | 13.47 | 22.06 |
| Country | 19.24 | 21.78 |
| Time | 12.98 | 24.79 |
Joint SFT.
To assess whether simultaneous optimization of spatial and temporal objectives reduces interference, we fine-tune a single model jointly on country and time prediction for 5 epochs, repeated across 3 runs to assess stability. Table 22 summarizes the averaged results. Joint training improves over the zero-shot baseline (country: 14.23% 15.72%; time: 20.27% 22.36%), but falls below the single-task peaks for both objectives (country: 19.24%; time: 24.79%). Time accuracy is more volatile across runs than country accuracy, consistent with the higher sensitivity of illumination cues to gradient noise. These results indicate that joint optimization under LoRA-based adaptation is limited by capacity constraints and competing gradient directions, suggesting that future work should explore gradient-disentangled or task-modular fine-tuning (schulzebuschoff2025finetuning; binz2025foundation).
| Epoch | Avg Country (%) | Avg Time (%) |
| 0 | 14.23 | 20.27 |
| 1 | 14.80 | 20.62 |
| 2 | 15.14 | 20.80 |
| 3 | 15.10 | 21.86 |
| 4 | 15.46 | 21.90 |
| 5 | 15.72 | 22.36 |
Future work should consider RL-based training strategies (chen2025sftrl) that allow more flexible task-specific reasoning rather than enforcing shared representations through SFT alone.
H.2 Joint Supervised Fine-Tuning
Single-task SFT treats country and time prediction as independent objectives, which leaves open the question of whether joint optimization over both tasks yields a more capable geo-temporal model. To investigate this, we fine-tune unsloth/Qwen2.5-VL-3B-Instruct simultaneously on country and time prediction under the same LoRA configuration and data split described in §H.1, extending training to five epochs. To assess result stability, we repeat the experiment across three independent runs (differing only in data ordering) and report per-run and averaged results.
Results.
Table 23 reports per-run and averaged accuracy for each epoch, with learning curves shown in Figures 10 and 11. Joint training improves over the zero-shot baseline on both tasks: averaged country accuracy rises from 14.23% to 15.72%, and averaged time accuracy rises from 20.27% to 22.36% by epoch 5. However, both values fall substantially below the corresponding single-task peaks (country: 19.24%; time: 24.79%), indicating a clear cost to joint optimization.
| Epoch | Run 1 | Run 2 | Run 3 | Average | ||||
| Country | Time | Country | Time | Country | Time | Country | Time | |
| 0 | 14.23 | 20.27 | 14.23 | 20.27 | 14.23 | 20.27 | 14.23 | 20.27 |
| 1 | 14.36 | 20.82 | 14.64 | 20.96 | 15.40 | 20.07 | 14.80 | 20.62 |
| 2 | 14.57 | 20.48 | 15.05 | 20.62 | 15.81 | 21.31 | 15.14 | 20.80 |
| 3 | 14.57 | 22.27 | 15.33 | 21.44 | 15.40 | 21.86 | 15.10 | 21.86 |
| 4 | 15.19 | 21.99 | 15.40 | 21.99 | 15.81 | 21.72 | 15.46 | 21.90 |
| 5 | 15.81 | 22.34 | 15.53 | 22.54 | 15.81 | 22.20 | 15.72 | 22.36 |
Cross-task evaluation.
Table 24 compares the zero-shot baseline, single-task SFT, and joint SFT on both metrics simultaneously. When fine-tuned on country alone, time accuracy drops from 22.06% to 21.78%; when fine-tuned on time alone, country accuracy drops from 13.47% to 12.98%. Joint SFT recovers both tasks above their degraded single-task cross-evaluation values but remains below single-task peaks.
| Training objective | Country Ac. (%) | Time Ac. (%) |
| Baseline (no SFT) | 13.47 | 22.06 |
| Single-task: Country | 19.24 | 21.78 |
| Single-task: Time | 12.98 | 24.79 |
| Joint SFT (avg, ep. 5) | 15.72 | 22.36 |
Analysis.
The performance gap between joint and single-task SFT reflects a structural conflict between the two objectives. Country prediction requires the model to be invariant to illumination: the same country must be recognized across daytime, nighttime, and seasonal conditions. Time prediction requires the opposite: sensitivity to illumination-dependent signals such as shadow length, sun elevation, sky luminance, and color temperature. Under LoRA-based adaptation, both objectives share the same low-rank update matrices. Satisfying their competing gradient directions simultaneously forces a compromise that suppresses task-specific specialization in both domains.
The asymmetry between country and time learning curves provides further evidence of this conflict. Country accuracy improves slowly and saturates early across all runs, indicating that the model struggles to acquire illumination-invariant place identity when the time objective simultaneously pulls adapter weights toward illumination sensitivity. Time accuracy, by contrast, shows a smoother upward trend but still falls below its single-task ceiling, reflecting that illumination-sensitive updates are partially dampened by the simultaneous country objective.
Implications and future directions.
These findings are consistent with known limitations of standard SFT on heterogeneous objectives (schulzebuschoff2025finetuning; binz2025foundation): enforcing a single shared representation tends to suppress the task-specific reasoning patterns that each objective individually supports. RL-based training approaches such as GRPO-style optimization (chen2025sftrl) could mitigate this by assigning per-task outcome rewards, allowing the model to develop more flexible reasoning strategies without being constrained to a single intermediate representation. Task-modular architectures with separate adapters per objective, or curriculum strategies that alternate between tasks at the gradient level, represent additional directions worth exploring in future work. Overall, joint SFT provides a useful multi-task baseline that confirms both tasks are learnable under shared supervision, while the gap relative to single-task training quantifies the interference cost and motivates more principled multi-task training designs.
@@ I.1 Limitations of Current Vision–Language Models
Limits of supervised fine-tuning. Supervised fine-tuning provides consistent gains for categorical geo-semantic prediction but yields smaller and less stable improvements for continuous temporal inference. Single-task SFT induces cross-task interference: country fine-tuning reduces time accuracy and vice versa, because the two tasks require competing visual sensitivities under shared LoRA parameters. Joint SFT partially mitigates this but falls below single-task peaks due to gradient conflict. SFT tends to reinforce frequent patterns aligned with structured labels, while struggling to induce robust reasoning over solar geometry and illumination physics, motivating architectural and training-level interventions beyond standard fine-tuning (chen2025sftrl; schulzebuschoff2025finetuning).
@@ I.2 Future Research Directions
Beyond fine-tuning: architectural and training innovations. While SFT improves baseline performance, our results indicate that deeper progress will require architectural changes and training paradigms that explicitly model time as part of world representation. Promising directions include hybrid perception–physics models, modular temporal reasoning components, and multi-task training that couples spatial perception with temporal prediction. Our cross-task interference findings also motivate RL-based approaches such as GRPO-style training (chen2025sftrl), which can support more flexible, task-aware reasoning strategies than SFT by operating on outcome-level rewards rather than enforcing shared intermediate representations. Multi-task SFT with explicit task routing or gradient-isolated adapters (schulzebuschoff2025finetuning; binz2025foundation) is another promising direction.
@@
Appendix J Human Performance and Comparison
J.1 Human Evaluation Protocol and Observed Patterns
We conducted a controlled human evaluation of TimeSpot using six participants divided into two cohorts: three undergraduate students (ages 20–22, generalists with no formal domain specialization) and three domain experts (Master’s students and senior PhD researchers with backgrounds in geomorphology, climatology, and geosciences). All participants completed the full benchmark under identical conditions, and their responses were evaluated using the same metrics as the VLMs.
The undergraduate cohort demonstrates moderate performance with strengths in coarse environmental recognition and daylight perception, but exhibits clear limitations in fine-grained localization and temporal precision. In contrast, the expert cohort consistently achieves high accuracy across both geo-spatial and temporal tasks, with particularly strong performance in climate classification, season inference, and time estimation.
Across both groups, two general patterns emerge: (i) human performance improves systematically with domain knowledge, especially for tasks requiring physical and environmental reasoning, and (ii) humans exhibit lower variance and more consistent cross-task performance compared to models, indicating a unified underlying reasoning framework rather than task-specific heuristics.
J.2 Detailed Comparison
The comparative evaluation between Vision-Language Models (VLMs) (Table 3) and human cohorts (Table 25) on TimeSpot reveals systematic and interpretable differences in how artificial and biological systems process geo-temporal information. While models demonstrate strong performance in discrete classification and large-scale memorization tasks, humans consistently outperform in domains requiring causal reasoning, physical intuition, and temporal inference. Below, we present a structured analysis across six key dimensions.
| Model / Subject | Geo-location Understanding | Temporal Understanding | ||||||||||
| Human Baselines | ||||||||||||
| Undergrad 1 (Generalist) | 81.57 | 46.28 | 68.56 | 75.04 | 12.57 | 32.16 | 2750.47 | 68.58 | 28.56 | 42.14 | 2:41 | 78.56 |
| Undergrad 2 (Traveler) | 84.06 | 52.18 | 71.26 | 78.57 | 11.26 | 28.57 | 2450.18 | 72.06 | 31.07 | 45.57 | 2:26 | 81.07 |
| Undergrad 3 (Less Traveled) | 77.08 | 39.56 | 64.07 | 73.58 | 14.86 | 38.46 | 3200.86 | 66.08 | 25.57 | 38.06 | 2:56 | 74.07 |
| Average (Undergrad) | 80.89 | 45.98 | 67.96 | 75.71 | 12.89 | 33.06 | 2800.49 | 68.89 | 28.39 | 41.92 | 2:41 | 77.89 |
| Expert 1 (Geomorphology) | 95.56 | 68.57 | 85.06 | 88.56 | 4.56 | 11.57 | 980.57 | 84.57 | 44.06 | 58.06 | 1:36 | 92.56 |
| Expert 2 (Climatology) | 92.07 | 64.06 | 89.57 | 86.07 | 5.87 | 14.27 | 1250.27 | 89.06 | 48.57 | 54.57 | 1:46 | 90.07 |
| Expert 3 (Field Geologist) | 94.57 | 71.08 | 84.56 | 87.56 | 5.06 | 10.86 | 890.47 | 86.07 | 45.57 | 61.07 | 1:26 | 94.07 |
| Average (Expert) | 94.06 | 67.89 | 86.39 | 87.39 | 5.16 | 12.22 | 1040.42 | 86.56 | 46.06 | 57.89 | 1:36 | 92.22 |
Geopolitical Recognition vs. Human Heuristics.
VLMs exhibit a clear advantage in country-level classification. The best-performing proprietary and reasoning models achieve accuracies exceeding , surpassing both undergraduate () and expert () human averages. This performance reflects the models’ ability to encode high-resolution statistical associations between visual patterns and geopolitical labels.
Humans, in contrast, rely on coarse semantic cues such as language, architectural style, or cultural artifacts. These cues are often insufficient for distinguishing between geographically or culturally similar regions. Consequently, human errors tend to cluster at geopolitical boundaries, where visual differences are subtle or non-existent.
This indicates that VLMs operate as high-capacity retrieval systems for geographically indexed visual features, whereas human reasoning is constrained by semantic abstraction and limited exposure to global diversity.
Environmental Semantics and Climate Reasoning. When the task shifts to climate zone and environment classification, the performance gap reverses. Human experts achieve accuracy in climate classification, significantly exceeding all model categories, where the best results plateau around . This difference highlights the importance of domain knowledge and multi-scale reasoning. Humans integrate vegetation structure, soil characteristics, and atmospheric conditions into coherent environmental interpretations. These features are continuous and interdependent, requiring abstraction beyond pixel-level correlations. VLMs, despite strong performance in environment type (up to ), struggle with climate categorization due to the absence of explicit supervision linking visual features to scientific taxonomies. This suggests that current models lack robust internal representations of ecological systems.
Metric Localization and Error Distribution. Latitude, longitude, and mean distance (MD) provide insight into spatial precision. Top VLMs achieve lower MD values (e.g., km) compared to experts ( km) and significantly outperform undergraduates ( km). However, this advantage must be interpreted cautiously. Model predictions are often anchored to high-confidence country-level guesses, leading to outputs near geographic centroids. This reduces average error but does not necessarily reflect true spatial reasoning. In contrast, human errors are less frequent but more dispersed, resulting in higher variance when incorrect. Notably, expert humans achieve latitude MAE () comparable to leading models (–), suggesting that humans employ physically grounded estimation strategies, such as solar elevation and vegetation gradients.
Temporal Inference and Solar Geometry. Temporal prediction represents the most significant failure mode for VLMs. Across all models, time-of-day accuracy remains below , with mean absolute errors consistently around hours. In contrast, undergraduates achieve MAE, and experts reduce this further to . This gap reflects the absence of explicit physical reasoning in current VLM architectures. Accurate time estimation requires interpreting shadow orientation, shadow length, atmospheric scattering, and human activity patterns. These cues are governed by deterministic physical laws, particularly solar geometry. Humans naturally integrate these signals into a coherent temporal estimate, while models treat them as weak statistical features. The resulting errors often manifest as systematic ambiguities, such as confusion between morning and late afternoon lighting conditions.
Seasonality and Biological Signals. Season classification further exposes differences in temporal reasoning. Human experts achieve accuracy, substantially outperforming all models (best ). Even undergraduates () remain competitive with top reasoning models. Season identification depends on phenological cues, including vegetation cycles, snow cover, and atmospheric clarity. These signals require longitudinal understanding of environmental change, which humans acquire through experience and education. Models, lacking explicit temporal grounding, rely on superficial cues such as color distributions. This leads to systematic weaknesses, particularly in transitional seasons where visual differences are subtle and context-dependent. Month prediction remains challenging for both groups due to the increased granularity (12 classes), but humans still maintain a consistent advantage, especially at the expert level ( vs. for top models).
Daylight Phase and Perceptual Coherence.
Daylight phase (DLP) accuracy provides a coarse measure of temporal understanding. Humans again demonstrate a strong advantage, with experts reaching and undergraduates , compared to a maximum of among models.
This task requires integrating global illumination characteristics, including color temperature, shadow softness, and sky luminance gradients. These features are perceptually salient to humans but remain weakly encoded in current VLM representations.
The discrepancy suggests that human perception is inherently calibrated to natural lighting conditions, enabling robust classification even under ambiguous scenarios.
The results collectively indicate a fundamental distinction between two modes of intelligence:
VLMs: High-dimensional statistical systems optimized for pattern recognition and large-scale memorization. Their strengths lie in discrete classification tasks with strong visual-textual correlations.
Humans: Physically grounded reasoning systems capable of integrating spatial, temporal, and environmental cues into coherent world models.
VLMs excel in tasks that resemble retrieval from a global visual database, such as country recognition. However, they lack the causal and geometric reasoning required for accurate temporal inference and environmental understanding. In contrast, humans interpret images as projections of dynamic physical processes. This enables robust performance in tasks involving time, climate, and environmental context, even with limited exposure to specific locations. Bridging this gap will require future models to incorporate explicit representations of physical laws, temporal dynamics, and environmental systems, moving beyond static visual pattern matching toward integrated spatio-temporal reasoning.
Appendix K Examples from TimeSpot
@@
Appendix L Examples of JSON record with Response Analysis